# SCATTER Common Envelope Formalism for Triples

### v1

@@ Abstract
Many stars are components of triple-star systems, or of higher-order multiples. In such systems mass transfer is common, and when the transfer is dynamically unstable, a common envelope forms. As such, it is important to be able to compute the post–common-envelope orbital separations among the various stars comprising the system, and to determine whether the common envelope induces mergers and/or makes later mergers inevitable. In this paper we compute the results of common-envelope evolution for triples. We employ the SCATTER formalism, a new approach to the computation of post–common-envelope separations. This work has applications to gravitational mergers, Type Ia supernovae, and a broad range of highly energetic phenomena.
@@
pubyear: 2025pagerange: SCATTER Common Envelope Formalism for TriplesB
@@ 1 Introduction
A large fraction of massive stars are born in triples (moe2017mind; duchene2013stellar; raghavan2010survey; TokovininTripleOrigin; tokovinin2014binaries). Stars in many triples become close enough to each other that the system will pass through at least one stage of mass transfer or mass loss. Of particular interest are scenarios leading to the merger of two or more of a triple’s stars. These scenarios generally require that the triple pass through at least one phase during which its components spiral closer to each other within a gaseous envelope that has been stripped from one of the companion stars. The envelope is called the common envelope (CE).
The CE phase of higher-order multiples is important because each CE phase can lead to mergers and/or to smaller orbits. Additionally, subsequent evolution of CE end states can lead to further mass transfer, mergers, and/or additional CE episodes. The CE process and its end states yield high-mass objects, alone or in multiple systems, that can eventually either exchange mass or merge. Whether mergers occur within the CE or afterwards, they can create extreme luminosities in the electromagnetic and/or gravitational-wave regimes. Some mergers produce Type Ia supernovae (SNe Ia), or accretion-induced collapse to an neutron star (NS) or black hole (BH). Triple-star CEs are expected to be common because massive stars have a high probability of starting as triple systems, and triples may be needed to provide pathways to certain intriguing end states, such as massive BHs. Furthermore, triples and higher-order multiples are also formed through dynamical interactions in dense environments (rasio1994binary; van2007formation; perets2012triple; leigh2011analytic; antonini2016black; martinez2020black; offner2022origin).
@@ 1.1 The CE in Binary Evolution
Common envelopes play important roles in binary evolution. The CE was first introduced as a way to form cataclysmic variables (pac76), binaries in which a white dwarf (WD) accretes mass from a close low-mass companion. Despite decades of study, there are still many open questions about CEs. Some present-day studies (passy2011simulating; ricker2012amr; ohlmann2015hydrodynamic; trani2022revisiting) focus on the complex task of simulating the evolution of individual systems (see ivanova2020common, Ch. 4, and ropke2023simulations for reviews). The most frequent use of the CE is in the context of population-synthesis studies. Population synthesis starts with a population of stellar systems and evolves each system to determine the number and characteristics of interesting end states, such as gravitational-wave mergers or SNe Ia. The evolutions are not the kind of detailed calculations one would do to predict the physical characteristics of each system over time. They are designed, however, to capture the significant features of stellar and binary evolution in a parameterized way, by using input from both theory and observations. While the end states predicted by the calculations may not be reliable guides to the characteristics of a particular individual system, the hope is that, on average, the ensemble of outcomes will represent realistic predictions for real populations of stars. Because a large fraction of “interesting” end-state systems pass through one or more intermediate states with a CE, population-synthesis calculations must be designed to map the state at the beginning of the CE to a state that has a good chance of realistically representing the final state of that or of a similar system.
In one set of CE formulations (the $\alpha$” formulation), the basic underlying principle is conservation of energy (van1976late; webbink1984; livio1988common). The energy needed to disperse the CE is provided at the expense of gravitational binding energy. While conservation of energy is a powerful principle, the nature and relative contributions of the sources and sinks of energy are many and varied, including recombination/ionization energy and shocks. An alternative approach is found in the $\gamma$” formalism, which applies conservation of angular momentum to the binary system as a whole (nelemans2000reconstructing; nelemans2005reconstructing).
@@
Here we apply the SCATTER formalism (DiStefano_2023), which is also based on the conservation of angular momentum. It allows each stellar component of the system to exchange angular momentum with the envelope, whether the system is a binary, triple, or higher-order multiple. Like the other formalisms, it maps an initial, pre-CE state to a final, post-CE state via an analytic formula. It includes, however, parameters that can be modeled as functions of mass ratios that vary from system to system. The choice of SCATTER parameter functions for binaries has been calibrated based on a nearly complete set of post–common-envelope binaries containing at least one WD (kruckow2021catalog). We therefore know that the version of SCATTER presented here is capable of describing a wide range of WD-containing post-CE systems.
The formalism provides enough freedom to incorporate data appropriate to any set of binaries or triples, including those that lead to BH–BH or BH–NS mergers. We do not know whether conducting a study of the post-CE states of such systems would produce the same set of parameterized functions derived for WD-containing systems. Such a comprehensive study is not yet possible, given the relatively small samples presently available for these massive systems. We therefore use parameter values derived from systems of lower mass and plan to incorporate new parameter values if needed, as more data become available.
@@ 1.2 The CE for Higher-Order Multiples
Adding a third body makes the problem more complex. There are now three potentially independent separations to consider: $a_{12},a_{13},$ and $a_{23}$. In principle, if we choose to consider the triple as a whole, we only have a single equation for total angular momentum, which alone is not enough to allow us to map the initial and final values of each of these three separations. Furthermore, it may be difficult to observationally measure the values of these separations either before or after the CE, especially if the separations are comparable to each other and the system is potentially chaotic. Even stable triples can exhibit a range of interesting behaviors linked to the exchange of angular momentum. For example, the motion of the outer star can alter the eccentricity and orientation of the inner binary (Kozai1962; LIDOV1962719; see Naoz2013 for a review).
@@
Furthermore, triples that respect a spatial hierarchy can be dynamically stable (Szebehely1967; harrington1968dynamical; eggleton1995empirical). Consider an inner binary composed of Stars 1 and 2, with masses $M_{1}$ and $M_{2}$, respectively. We denote their orbital separation by $a_{\rm in}$. If Star 3, with mass $M_{3}$, orbits the center of mass of the inner binary in a wide enough orbit, with semimajor axis $a_{\mathrm{out}}$, the system can be considered to be a hierarchical triple that is dynamically stable. Furthermore, for hierarchical triples in a CE, the inner and outer orbits can be considered separately. This is because the transfer of angular momentum to the envelope is more significant than transfers due to the direct interactions of the stars. Thus, rather than trying to solve for three unique orbital separations solely from the triple’s total angular momentum, we instead subdivide the triple into separate binaries to which we can then apply SCATTER to solve for the change in orbital separation.
@@ 2.1 Overview of SCATTER
The underlying assumption of SCATTER is that, during the CE, changes in the configuration of stars occur through interactions between each star and the matter in the CE. Thus, changes in the $N$-body system’s angular momentum are caused by the interaction of each body with the CE. Changes in inter-star distances are associated with the changes in angular momentum. Here, we summarize the results of the calculations for binaries. Details can be found in DiStefano_2023.
@@ 2.2 SCATTER for Binaries
In order to find appropriate values of $\eta$, DiStefano_2023 used the equation for $a(f)/a(0)$ to map the pre-CE state to the post-CE state of 112 post-CE binaries from kruckow2021catalog. This allowed a functional form for $\eta$ to be derived as
@@
In DiStefano_2023 a range of $\delta$ values was considered. In order to avoid a large dependence on the value of $\eta$ needed to achieve a specific end state, we found $\delta=3$ to be a good choice. This value is also associated with a clear physical interpretation, as $(a\times f(q))^{3}$ (where $a$ is the orbital separation) represents a volume.
@@ 2.3 Goals of the CE Calculations
We note that the SCATTER formalism often predicts less shrinkage than the $\alpha$ or $\gamma$ formalism (DiStefano_2023). SCATTER even predicts expansion in cases for which mass loss is more important than angular-momentum loss. Nevertheless, the amount of shrinkage predicted by SCATTER is enough to produce both prompt and delayed mergers for a wide range of systems.
@@ 3.2 Orbital Stability and Hierarchical Triples
Stability conditions have been derived, allowing us to identify stable triples. We use the stability criterion of Mardling1999 for prograde coplanar triples, which states that
@@
where $a_{\mathrm{out}}$ and $a_{\mathrm{in}}$ are the semimajor axes of the outer binary and the inner binary, respectively. We use $q_{\mathrm{out}}=\frac{M_{3}(0)}{M_{1}(0)+M_{2}(0)}$. The eccentricity of the outer orbit is $e_{\mathrm{out}}$. Thus, to ensure stability, the value of $a_{\mathrm{out}}$ must be greater than $G\times a_{\mathrm{in}}$. Because the stability condition requires that $a_{\mathrm{out}}$ be significantly larger than $a_{\mathrm{in}}$, it is appropriate to refer to systems satisfying the stability condition as hierarchical. When the stability condition is satisfied, calculations that treat the inner and outer binaries separately are likely to produce reasonable approximations to the results achieved by real triples. In this paper we set the eccentricity to zero.
@@ 3.3 The Roche Lobe
The concept of the Roche lobe (RL) emerged from the study of close binaries. When a star fills its RL, it is poised to transfer mass to a co-rotating companion. The RL formalism is modified for triples (dynamicalrl). The effects of the modification include a moving $L_{1}$ point. Changes from the standard formalism become more pronounced when the separations between the stars in the triple are of similar sizes. The conditions for the equivalent of RL filling are not significantly affected, especially for hierarchical triples, and we will use them here. The RL-filling star (i.e., the star that contributes its envelope) may be a member of either the inner or outer binary.
@@ 3.5 Post-CE States: Binaries
    Is there a merger during the CE? A merger during the CE will occur if the inspiral time of the binary due to drag is shorter than the time needed to expel the envelope. For the case of an input–output style formalism such as SCATTER we select binaries in which the time to merger post CE is less than $(10^{4}$$10^{5})$ yr (typical dispersion time for CE). Thus we assume that a merger takes place before the CE has fully dispersed, even if no envelope material remains near the site of the merger. When one component of the binary is an extended star, then the merger criteria are related to the stellar radius. The criteria depend on the nature of the star, but mergers are generally expected when the distance of closest approach is less than a few stellar radii. To be specific, we will say that a merger occurs when the extended star would overfill its RL at closest approach.
@@
    If there is a merger, what is the result? If the two components are compact objects, gravitational and possibly electromagnetic energy and a variety of particles will be released upon merger. The result of a BH–BH merger is a BH, as is the result of a BH–NS merger. For a BH–WD or a NS–WD system, the WD will be tidally disrupted and a significant fraction of the mass will be accreted by the WD’s companion. In general, the merger of two NSs will produce an object more massive than the maximum NS mass, potentially leading to an accretion-induced collapse (AIC) to a BH (bernuzzi2020neutron). The merger of two WDs can produce a SN Ia. Alternatively, WD mergers can lead to AIC to a NS, or simply to fast-rotating WDs or RCB stars (yungelson2017merging; dan2014structure). When one component is a main-sequence star or giant, mergers can yield fast-rotating giants, luminous red-nova–like transients, or (in WD/NS–giant cases) Thorne–Żytkow objects (ablimit2022stellar).
@@
    If there is no prompt merger, will the post-CE binary interact later? As above, the answers to this question depend on the natures of the two objects that remain in orbit with each other. For compact objects we need only ask if the time to gravitational merger is smaller than a Hubble time. If so, they will merge, producing the end states described above, but at a later time. If one of the stars is extended, then we must ask whether it will fill its RL in a Hubble time, thus leading to a second epoch of mass transfer that, depending on the relative masses and state of evolution of the RL-filling star, could lead to either a second CE or else an epoch of stable mass transfer. If there is no future RL filling, mass may be transferred through winds, depending on the mass of the extended star and the distance between it and its compact companion.
@@ 3.6 Post-CE End States: Triples
  • A triple will result in cases in which no merger occurs. If the stability limit is respected and the triple is dynamically stable, then the triple end state may be long lived. If the orbits are close enough that angular momentum is drained through the emission of gravitational radiation during the system’s evolution, then the effects of general relativity must be taken into account to assess stability. In either the relativistic or non-relativistic case, the triple may be chaotic. In this case, the system’s orbital evolution could lead to a triple-star merger or a double-star merger. Whether or not there is a merger, one of the three stars can be ejected at high velocity. If chaos reigns, then we can give only a probabilistic estimate of the outcomes (stone2019statistical; toonen2022stellar).

### v2

@@ Abstract
Many stars are components of triple-star systems, or of higher-order multiples. In such systems mass transfer is common, and when the transfer is dynamically unstable, a common envelope forms. As such, it is important to be able to compute the post–common-envelope orbital separations among the various stars comprising the system, and to determine whether the common envelope induces mergers or makes later mergers inevitable. In this paper we compute the results of common-envelope evolution for triples. We employ the SCATTER formalism, a new approach to the computation of post–common-envelope separations. This work has applications to gravitational-wave mergers, Type Ia supernovae, and a broad range of other highly energetic phenomena.
@@
pubyear: 2025pagerange: SCATTER Common Envelope Formalism for TriplesSCATTER Common Envelope Formalism for Triples
@@ 1 Introduction
A large fraction of massive stars are born in triples (Moe & Di Stefano, 2017; Duchêne & Kraus, 2013; Raghavan et al., 2010; Tokovinin, 2008, 2014). Stars in many triples become close enough to each other that the system will pass through at least one stage of mass transfer or mass loss. Of particular interest are scenarios leading to the merger of two or more of a triple’s stars. These scenarios generally require that the triple pass through at least one phase during which its components spiral closer to each other within a gaseous envelope that has been stripped from one of the companion stars. The envelope is called the common envelope (CE).
The CE phase of higher-order multiples is important because each CE phase can lead to mergers and/or to smaller orbits. Additionally, subsequent evolution of CE end states can lead to further mass transfer, mergers, and/or additional CE episodes. The CE process and its end states yield high-mass objects, alone or in multiple systems, that can eventually either exchange mass or merge. Whether mergers occur within the CE or afterwards, they can create extreme luminosities in the electromagnetic and/or gravitational-wave regimes. Some mergers produce Type Ia supernovae (SNe Ia), or accretion-induced collapse to an neutron star (NS) or black hole (BH). Triple-star CEs are expected to be common because massive stars have a high probability of starting as triple systems (Moe & Di Stefano, 2017; Tokovinin, 2018), and triples may be needed to provide pathways to certain intriguing end states, such as binary BH mergers (Stegmann et al., 2025) and massive BHs [^1] This is because triples provide additional channels for interaction, beyond those available in binaries.. Furthermore, triples and higher-order multiples are also formed through dynamical interactions in dense environments (Rasio et al., 1994; Van Den Berk et al., 2007; Perets & Kratter, 2012; Leigh & Sills, 2011; Antonini et al., 2016; Martinez et al., 2020; Offner et al., 2022).
@@ 1.1 The CE in Binary Evolution
Common envelopes play important roles in binary evolution. The CE was first introduced as a way to form cataclysmic variables (Paczynski, 1976), binaries in which a white dwarf (WD) accretes mass from a close low-mass companion. Despite decades of study, there are still many open questions about CEs. Some present-day studies (Passy et al., 2011; Ricker & Taam, 2012; Ohlmann et al., 2015; Trani et al., 2022) focus on the complex task of simulating the evolution of individual systems (see Ivanova et al. (2020), Ch. 4, and Röpke & De Marco (2023) for reviews). The most frequent use of the CE is in the context of population-synthesis studies. Population synthesis starts with a population of stellar systems and evolves each system to determine the number and characteristics of interesting end states, such as gravitational-wave mergers or SNe Ia. The evolutions are not the kind of detailed calculations one would do to predict the physical characteristics of each system over time. They are designed, however, to capture the significant features of stellar and binary evolution in a parameterized way, by using input from both theory and observations. While the end states predicted by the calculations may not be reliable guides to the characteristics of a particular individual system, the hope is that, on average, the ensemble of outcomes will represent realistic predictions for real populations of stars. Because a large fraction of “interesting” end-state systems pass through one or more intermediate states with a CE, population-synthesis calculations must be designed to map the state at the beginning of the CE to a state that has a good chance of realistically representing the final state of that or of a similar system. It has been noted that, both for the production of Type Ia supernovae through the mergers of WDs and for the production of compact-object mergers that involve BHs or NSs, the CE evolution plays a key role (Ruiter et al., 2009; Olejak et al., 2021).
In one set of CE formulations (the $\alpha$” formulation), the basic underlying principle is conservation of energy (Van Den Heuvel, 1976; Webbink, 1984; Livio & Soker, 1988). The energy needed to disperse the CE is provided at the expense of gravitational binding energy. While conservation of energy is a powerful principle, the nature and relative contributions of the sources and sinks of energy are many and varied, including recombination/ionization energy and shocks. An alternative approach is found in the $\gamma$” formalism, which applies conservation of angular momentum to the binary system as a whole (Nelemans et al., 2000; Nelemans & Tout, 2005).
@@
Here we apply the SCATTER formalism (Di Stefano et al., 2023), which is also based on the conservation of angular momentum. It allows each stellar component of the system to exchange angular momentum with the envelope, whether the system is a binary, triple, or higher-order multiple. Like the other formalisms, it maps an initial, pre-CE state to a final, post-CE state via an analytic formula. It includes, however, parameters that can be modeled as functions of mass ratios that vary from system to system. The choice of SCATTER parameter functions for binaries has been calibrated based on a nearly complete set of post–common-envelope binaries containing at least one WD (Kruckow et al., 2021). We therefore know that the version of SCATTER presented here is capable of describing a wide range of WD-containing post-CE systems.
The formalism provides enough freedom to incorporate data appropriate to any set of binaries or triples, including those that lead to BH–BH, NS–NS, or BH–NS mergers. We do not know whether conducting a study of the post-CE states of such systems would produce the same set of parameterized functions derived for WD-containing systems. Such a comprehensive study is not yet possible, given the relatively small samples presently available for these massive systems. We therefore use parameter values derived from systems of lower mass and plan to incorporate new parameter values if needed, as more data become available.
@@ 1.2 The CE for Higher-Order Multiples
Adding a third body makes the problem more complex. There are now three potentially independent separations to consider: $a_{12},a_{13},$ and $a_{23}$. In principle, if we choose to consider the triple as a whole, we only have a single equation for total angular momentum, which alone is not enough to allow us to map the initial and final values of each of these three separations. Furthermore, it may be difficult to observationally measure the values of these separations either before or after the CE, especially if the separations are comparable to each other and the system is potentially chaotic. Even stable triples can exhibit a range of interesting behaviors linked to the exchange of angular momentum. For example, the motion of the outer star can alter the eccentricity and orientation of the inner binary (Kozai 1962; Lidov 1962; see Naoz et al. 2013 for a review).
@@
Furthermore, triples that respect a spatial hierarchy can be dynamically stable (Szebehely & Peters, 1967; Harrington, 1968; Eggleton & Kiseleva, 1995). Consider an inner binary composed of Stars 1 and 2, with masses $M_{1}$ and $M_{2}$, respectively. We denote their orbital separation by $a_{\rm in}$. If Star 3, with mass $M_{3}$, orbits the center of mass of the inner binary in a wide enough orbit, with semimajor axis $a_{\mathrm{out}}$, the system can be considered to be a hierarchical triple that is dynamically stable. Furthermore, for hierarchical triples in a CE, the inner and outer orbits can be considered separately. This is because the transfer of angular momentum to the envelope is more significant than transfers due to the direct interactions of the stars. Thus, rather than trying to solve for three unique orbital separations solely from the triple’s total angular momentum, we instead subdivide the triple into separate binaries to which we can then apply SCATTER to solve for the change in orbital separation.
@@ 2.1 Overview of SCATTER
The underlying assumption of SCATTER is that, during the CE, changes in the configuration of stars occur through interactions between each star and the matter in the CE. Thus, changes in the $N$-body system’s angular momentum are caused by the interaction of each body with the CE. Changes in inter-star distances are associated with the changes in angular momentum. Here, we summarize the results of the calculations for binaries. Details can be found in Di Stefano et al. (2023).
@@ 2.2 SCATTER for Binaries
In order to find appropriate values of $\eta$, Di Stefano et al. (2023) used the equation for $a(f)/a(0)$ to map the pre-CE state to the post-CE state of 112 post-CE binaries from Kruckow et al. (2021). This allowed a functional form for $\eta$ to be derived as
@@
In Di Stefano et al. (2023) a range of $\delta$ values was considered. In order to avoid a large dependence on the value of $\eta$ needed to achieve a specific end state, we found $\delta=3$ to be a good choice. This value is also associated with a clear physical interpretation, as $(a\times f(q))^{3}$ (where $a$ is the orbital separation) represents a volume.
@@ 2.3 Goals of the CE Calculations
We note that the SCATTER formalism often predicts less shrinkage than the $\alpha$ or $\gamma$ formalism (Di Stefano et al., 2023). SCATTER even predicts expansion in cases for which mass loss is more important than angular-momentum loss. Nevertheless, the amount of shrinkage predicted by SCATTER is enough to produce both prompt and delayed mergers for a wide range of systems.
@@ 3.2 Orbital Stability and Hierarchical Triples
Stability conditions have been derived, allowing us to identify stable triples. We use the stability criterion of Mardling & Aarseth (1999) for prograde coplanar triples, which states that
@@
where $a_{\mathrm{out}}$ and $a_{\mathrm{in}}$ are the semimajor axes of the outer binary and the inner binary, respectively. We use $q_{\mathrm{out}}=\frac{M_{3}(0)}{M_{1}(0)+M_{2}(0)}$. The eccentricity of the outer orbit is $e_{\mathrm{out}}$. Thus, to ensure stability, the value of $a_{\mathrm{out}}$ must be greater than $G\times a_{\mathrm{in}}$. Because the stability condition requires that $a_{\mathrm{out}}$ be significantly larger than $a_{\mathrm{in}}$, it is appropriate to refer to systems satisfying the stability condition as hierarchical. When the stability condition is satisfied, calculations that treat the inner and outer binaries separately are likely to produce reasonable approximations to the results achieved by real triples. In this paper we set the eccentricity to zero. The form of Equation 6 assumes co-planarity, although not all systems undergoing a CE phase are co-planar. It has been shown, however, that the form of the stability criterion is not a strong function of orbital inclination (Eggleton & Kiseleva, 1995).
@@ 3.3 The Roche Lobe
The concept of the Roche lobe (RL) emerged from the study of close binaries. When a star fills its RL, it is poised to transfer mass to a co-rotating companion. The RL formalism is modified for triples (Di Stefano, 2019). The effects of the modification include a moving $L_{1}$ point. Changes from the standard formalism become more pronounced when the separations between the stars in the triple are of similar sizes. The conditions for the equivalent of RL filling are not significantly affected, especially for hierarchical triples, and we will use them here. The RL-filling star (i.e., the star that contributes its envelope) may be a member of either the inner or outer binary.
@@ 3.5 Post-CE States: Binaries
    Is there a merger during the CE? A merger during the CE will occur if the inspiral time of the binary due to drag is shorter than the time needed to expel the envelope. For the case of an input–output style formalism such as SCATTER we select binaries in which the time to merger post CE is less than $(10^{4}$$10^{5})$ yr (typical dispersion time for CE). Thus we assume that a merger takes place before the CE has fully dispersed, even if no envelope material remains near the site of the merger. When one component of the binary is an extended star, then the merger criteria are related to the stellar radius. The criteria depend on the nature of the star, but mergers are generally expected when the distance of closest approach is less than a few stellar radii. To be specific, we will say that a merger occurs when the extended star would overfill its RL at closest approach. For a binary whose only components are BHs or NS, we say that a merger will occur during the CE if the final orbital separation yields a time to merger shorter than $10^{4}$ years. This value is a proxy for the lifetime of the CE; although our choice is somewhat arbitrary, the value can be tailored to suit specific types of binaries as needed.
@@
    If there is a merger, what is the result? If the two components are compact objects, gravitational and possibly electromagnetic energy and a variety of particles will be released upon merger. The result of a BH–BH merger is a BH, as is the result of a BH–NS merger. For a BH–WD or a NS–WD system, the WD will be tidally disrupted and a significant fraction of the mass will be accreted by the WD’s companion. In general, the merger of two NSs will produce an object more massive than the maximum NS mass, potentially leading to an accretion-induced collapse (AIC) to a BH (Bernuzzi, 2020). The merger of two WDs can produce a SN Ia. Alternatively, WD mergers can lead to AIC to a NS, or simply to fast-rotating WDs or RCB stars (Yungelson & Kuranov, 2017; Dan et al., 2014). When one component is a main-sequence star or giant, mergers can yield fast-rotating giants, luminous red-nova–like transients, or (in WD/NS–giant cases) Thorne–Żytkow objects (Ablimit et al., 2022).
@@
    If there is no prompt merger, will the post-CE binary interact later? As above, the answers to this question depend on the natures of the two objects that remain in orbit with each other. For compact objects we need only ask if the time to gravitational-wave merger is smaller than a Hubble time. If so, they will merge, producing the end states described above, but at a later time. If one of the stars is extended, then we must ask whether it will fill its RL in a Hubble time, thus leading to a second epoch of mass transfer that, depending on the relative masses and state of evolution of the RL-filling star, could lead to either a second CE or else an epoch of stable mass transfer. If there is no future RL filling, mass may be transferred through winds, depending on the mass of the extended star and the distance between it and its compact companion.
@@ 3.6 Post-CE End States: Triples
  • A triple will result in cases in which no merger occurs. If the stability limit is respected and the triple is dynamically stable, then the triple end state may be long lived. If the orbits are close enough that angular momentum is drained through the emission of gravitational radiation during the system’s evolution, then the effects of general relativity must be taken into account to assess stability. In either the relativistic or non-relativistic case, the triple may be chaotic. In this case, the system’s orbital evolution could lead to a triple-star merger or a double-star merger. Whether or not there is a merger, one of the three stars can be ejected at high velocity. If chaos reigns, then we can give only a probabilistic estimate of the outcomes (Stone & Leigh, 2019; Toonen et al., 2022).
Note that we are considering that interval in the evolution during which the interactions of the CE with the multiple-star system dominate. There are cases, particularly when two components of the system are NSs or BHs, when gravitational effects can become more important, especially during close approaches. For example, one may speculate that depending on the natures and relative masses of the triple’s components at the time that two of the stars merge, the merger could cause a significant recoil, a gravitational-wave “kick” with enough energy and momentum to potentially expel the remaining companion if the efficiency of energy and momentum transfer is high enough (see also Naoz et al. 2025). This possibility should be included in calculations when the natures of the objects undergoing potential mergers are known to include any combination of NSs and BHs. The SCATTER formalism (or any of the other CE formalisms) can be employed up until the time when general-relativistic effects become as significant as the CE in changing orbital separations.